OpenStudy (anonymous):

How many mph is a car traveling if the wheels have a diameter of 3ft and are turning 600 times a minute

OpenStudy (lucaz):

circunference = pi*diameter = 3*pi ft 600 times a minute = 600*3pi ft / min you have to convert this into miles / hour

OpenStudy (lucaz):